AnimeGhoul8863:

Given the equation 4x + 7 = 3(2x − 5), solve for the variable. Explain each step and justify your process. Sam solved a similar equation below. Is Sam's solution correct? Explain why or why not. 5x − 1 = 2(x + 4) 5x − 1 = 2x + 8 7x − 1 = 8 7x = 7 x = 1

AnimeGhoul8863:

@Shadow

AnimeGhoul8863:

@TheSmartOne

TheSmartOne:

Do you see something wrong with his work

AnimeGhoul8863:

7x − 1 = 8 7x = 7 x = 1

TheSmartOne:

Good catch, I didn't see that one! Yeah they didn't add the 1 properly on both sides But look at them line before that too There's one more mistake they started off with Hint: 2x

AnimeGhoul8863:

2(x + 4) 2x + 8

TheSmartOne:

No, when they're supposed to subtract the 2x

AnimeGhoul8863:

i just noticed were doing the second part

TheSmartOne:

Oh, totally forgot there was a question before that

TheSmartOne:

Do we really have to justify the process and explain our steps

AnimeGhoul8863:

yes i think so cause it says to

TheSmartOne:

That's too much work, but solve it out first Do the distributing and isolating the variables

AnimeGhoul8863:

srry im trying to think how i do this

TheSmartOne:

Not much to think about 3(2x-5) Distribute the 3

AnimeGhoul8863:

3(2x-5) turns into 6x-5?

TheSmartOne:

Uh no a(b+c) = ab + ac

TheSmartOne:

Likewise a(b-c) = ab - ac

AnimeGhoul8863:

so 3*2-5*3

TheSmartOne:

You can't forget about the x, which is a variable

AnimeGhoul8863:

so 3*2x-5*3

AnimeGhoul8863:

u there

TheSmartOne:

Yes, simplify that

BenLindquist:

i know where he messed up and what should have been done. I don't know what you are on with tso

AnimeGhoul8863:

ok ben and TSO if im correct it =3

TheSmartOne:

All yours then Ben

BenLindquist:

i know how but i can't without giving the answer

BenLindquist:

i know how but i can't without giving the answer

TheSmartOne:

Also there were two questions

BenLindquist:

yeah i have both. its on my doc

AnimeGhoul8863:

so ben ami right is it 3

BenLindquist:

yes

AnimeGhoul8863:

so Part A: Step 1:4x + 7 = 3(2x − 5) Step 2: 3(2x-5) Distribute the 3 Step 3: 3*2x-5*3 Step 4: x=3 simplify

AnimeGhoul8863:

am i correct so far

AnimeGhoul8863:

what?????????????????????

AnimeGhoul8863:

the answer is 3 for both equations

AnimeGhoul8863:

im sooooooo confused

BenLindquist:

Question 1

BenLindquist:

A. Given the equation 4x + 7 = 3(2x − 5), solve for the variable.  4x + 7 = 3(2x − 5) 4x + 7 = 6x - 15 4x - 6x = -15 - 7 -2x = -22 x = 11

BenLindquist:

Question 2

BenLindquist:

5x − 1 = 2(x + 4) 5x − 1 = 2x + 8 5x − 2x = 8 + 1 3x = 9 x = 3 Sam messed up at line 2

AnimeGhoul8863:

thx but im still confused TSO what were u trying to show me cause what u had is soooooooo different than what ben has

TheSmartOne:

I was going step by step, he just threw all the steps